# VLEU: a Method for Automatic Evaluation for Generalizability of Text-to-Image Models

### v1

@@ Abstract
Progress in Text-to-Image (T2I) models has significantly advanced the generation of images from textual descriptions. However, there is a lack of metrics to evaluate the models’ ability to handle a diverse and extensive range of textual prompts—a key aspect of generalizability. To fill this gap, we propose the VLEU (Visual Language Evaluation Understudy) metric. VLEU leverages the power of the Large Language Model (LLM) to sample from the visual text domain, which we define as the entire set of potential inputs for the T2I task, to generate a wide variety of visual text. The images generated by T2I models from these prompts are then assessed for their alignment with the input text using the CLIP model. VLEU quantitatively measures a model’s generalizability by computing the Kullback-Leibler (KL) divergence between the visual text marginal distribution and the conditional distribution over the images generated by the model. This measure serves as a quantitative metric for comparing the generalizability of T2I models, and also provides a valuable metric for assessing improvements during the finetuning process. Our experimental results demonstrate VLEU’s effectiveness in evaluating the generalizability of various T2I models, positioning it as an essential metric for future research and development in image synthesis from text prompts. Our code and data will be publicly available at https://github.com.
@@ 3 VLEU
Figure 2: Generalizability in T2I Models: A Comparative Visualization. Model G demonstrates strong generalization, successfully producing diverse images from three distinct prompts. Conversely, $\textbf{G}^{\prime}$ shows limitations, overfitting to the prompt "A yellow dog" and failing to generalize to other inputs, resulting in similar outputs. Our proposed VLEU metric aims to quantify this observation.
@@ 4.1 Effectiveness Analysis
We selected the finetuning process of T2I models on specific datasets, which is often considered detrimental to the models’ generalizability, and tested the changes of IS (Salimans et al., 2016), FID (Heusel et al., 2017), CLIP score and VLEU during finetuning. [^3] For the computation of IS and FID, we utilized toshas/torch-fidelity. Specifically, we finetuned SD 1.5 on the DreamBooth dataset and sampled 25 prompts related to each subset using GPT 3.5. For FID calculation, we treated the images in the dataset as the real image distribution. For CLIP score, we simply calculate the average clip similarity of all prompts and corresponding image.

### v2

@@ Abstract
Progress in Text-to-Image (T2I) models has significantly advanced the generation of images from textual descriptions. Existing metrics, such as CLIP, effectively measure the semantic alignment between single prompts and their corresponding images. However, they fall short in evaluating a model’s ability to generalize across a broad spectrum of textual inputs. To address this gap, we propose the VLEU (Visual Language Evaluation Understudy) metric. VLEU leverages the power of Large Language Models (LLMs) to sample from the visual text domain, encompassing the entire range of potential inputs for the T2I task, to generate a wide variety of visual text. The images generated by T2I models from these prompts are then assessed for their alignment with the input text using the CLIP model. VLEU quantitatively measures a model’s generalizability by computing the Kullback-Leibler (KL) divergence between the visual text marginal distribution and the conditional distribution over the images generated by the model. This provides a comprehensive metric for comparing the overall generalizability of T2I models, beyond single-prompt evaluations, and offers valuable insights during the finetuning process. Our experimental results demonstrate VLEU’s effectiveness in evaluating the generalizability of various T2I models, positioning it as an essential metric for future research and development in image synthesis from text prompts. Our code and data will be publicly available at https://github.com/mio7690/VLEU.
@@ 3 VLEU
Figure 2: Generalizability in T2I Models: A Comparative Visualization. The first model demonstrates strong generalization by successfully producing images that align with various input prompts. In contrast, the second model shows overfitting to the prompt "A white and yellow dog," leading to a failure to generalize to other inputs. As a result, its generated images are generally misaligned with the given prompts. Our proposed VLEU metric aims to quantify this observation.
@@ 4.1 Effectiveness Analysis
We selected the finetuning process of T2I models on specific datasets, which is often considered detrimental to the models’ generalizability, and tested the changes of IS (Salimans et al., 2016), FID (Heusel et al., 2017), CLIP score and VLEU during finetuning. [^3] For the computation of IS and FID, we utilized toshas/torch-fidelity. Specifically, we finetuned SD 1.5 on the DreamBooth dataset and sampled 25 prompts related to each subset using GPT 3.5. For FID calculation, we treated the images in the dataset as the real image distribution. Note that the FID results may be affected by the small size of the image set. For CLIP score, we simply calculate the average clip similarity of all prompts and corresponding image.
@@ 8 Ethical Considerations
9 Acknowledgement
This research work is partially supported by ITF (No. PRP/054/21FX), CUHK direct grant (No. 4055209), and CUHK Knowledge Transfer Project Fund (No. KPF23GWP20).
@@ Appendix C Hyperparameter Analysis
Appendix D Prompt Diversity
To investigate the diversity of prompts generated by different LLMs, we generated word clouds for GPT-3.5, GPT-4, LLaMA-2-7B, and LLaMA-2-13B based on 1000 unconstrained prompts sampled from each model. As depicted in Figure 7, the word clouds reveal notable differences in vocabulary richness and diversity among the prompts generated by these models.
(a) ChatGPT 3.5
(b) GPT-4
(c) LLaMA-2-7B
(d) LLaMA-2-13B
Figure 7: Word clouds of sampled prompts. Each word cloud is generated from 1000 prompts, which were sampled by different LLMs on unconstrained subjects.
From the word clouds, it is evident that prompts sampled from GPT-4 and GPT-3.5 exhibit a richer and more diverse vocabulary. In these word clouds, there is a greater variety of words, and no single word overly dominates, indicating a balanced and broad coverage of themes and subjects. On the other hand, the word clouds generated from LLaMA-2-7B and LLaMA-2-13B display a limited vocabulary with a few high-frequency words occupying a significant portion of the visual space. This suggests that these models produce less diverse prompts.
We observed that utilizing LLMs with better generalizability, such as GPT-4, as the Visual Text Sampler, resulted in higher VLEU scores compared to those obtained with LLMs with poorer generalizability, such as LLaMA-2-7B. This indicates that prompts from models like GPT-4 describe the images with more detailed and varied language, leading to a higher alignment with the generated images, thereby reflecting the model’s generalizability more accurately. In contrast, LLMs with weaker descriptive capabilities produce prompts with substantially lower semantic alignment with the generated images, increasing the KL divergence and yielding higher VLEU scores. The generated word clouds and observed correlation between prompt diversity and VLEU scores underscore the importance of diverse and rich prompt generation in evaluating T2I models’ generalizability.